### v3

# SpeechRole: A Large-Scale Dataset and Benchmark for Evaluating Speech Role-Playing Agents

Abstract
Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance. Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios. In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs). To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations. Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing. Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity. Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence. We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
Abstract
Recently, role-playing agents have emerged as a promising paradigm for achieving personalized interaction and emotional resonance. Existing research primarily focuses on the textual modality, neglecting the critical dimension of speech in realistic interactive scenarios. In particular, there is a lack of systematic evaluation for Speech Role-Playing Agents (SRPAs). To address this gap, we construct SpeechRole-Data, a large-scale, high-quality dataset that comprises 98 diverse roles and 112k speech-based single-turn and multi-turn conversations. Each role demonstrates distinct vocal characteristics, including timbre and prosody, thereby enabling more sophisticated speech role-playing. Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses SRPAs performance in key aspects such as fundamental interaction ability, speech expressiveness, and role-playing fidelity. Experimental results reveal the advantages and challenges of both cascaded and end-to-end speech role-playing agents in maintaining vocal style consistency and role coherence. We release all data, code, and baseline models to provide a solid foundation for speech-driven multimodal role-playing research and to foster further developments in this field.
[…]
1 Introduction
Recently, Role-Playing Agents (RPAs) have emerged as a promising paradigm for achieving personalized interactions and emotional resonance [8]. Simulated characters with distinct traits make interactions more engaging in applications such as digital assistants [43], educational tools [44], and interactive storytelling [36]. However, current research predominantly focuses on the text modality [23, 32, 36, 31], overlooking the crucial role of speech in authentic interactions.
Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [5]. Although preliminary studies have explored role-playing incorporating visual or speech modalities [11, 46], there is currently a lack of systematic frameworks for constructing and evaluating Speech Role-Playing Agents (SRPAs). SRPAs are systems capable of engaging in character-driven spoken dialogue, exhibiting consistent vocal traits, expressive prosody, and coherent persona behavior across interactions. In particular, large-scale, high-quality datasets and evaluation benchmarks for modeling and assessing character consistency, prosodic variation, and interaction capabilities remain scarce.
To address this gap, we propose SpeechRole, a unified framework for the development and evaluation of SRPAs (see Figure 1). We first introduce SpeechRole-Data, a large-scale dataset comprising 112k speech-based dialogues covering 98 diverse roles. Each role features distinctive timbre and prosodic characteristics, enabling expressive and role-consistent speech interactions.
Furthermore, we propose SpeechRole-Eval, a multidimensional evaluation benchmark that systematically assesses agents across three dimensions: interaction competence, speech expressiveness, and role consistency. This benchmark supports both cascaded and end-to-end architectures, facilitating comprehensive and comparable analyses across different paradigms.
Our experiments reveal complementary strengths and weaknesses between SRPA paradigms. Cascaded systems offer strong modular control and stable role identity across interactions, but may suffer from cross-module mismatches or limited expressiveness. In contrast, end-to-end models produce more coherent and natural speech, yet often struggle with speaker consistency and preserving character traits over extended dialogues. These trade-offs highlight the unique challenges in building effective SRPAs that balance interaction ability, speech expressiveness, and role-playing fidelity.
Overall, our main contributions are as follows:
1.
We construct SpeechRole-Data, a large-scale multilingual corpus covering 98 roles and 112k speech-based dialogues with rich prosodic and timbral variation.
2.
We develop SpeechRole-Eval, a comprehensive benchmark designed to evaluate SRPA capabilities across key dimensions including interaction ability, speech expressiveness, and role-playing fidelity.
3.
We systematically compare cascaded and end-to-end paradigms, revealing their strengths, limitations, and offering insights for future SRPA development.
2 Related Work
Role-Playing Agents Across Modalities.
Recent research has increasingly explored the role-playing capabilities of Large Language Models (LLMs), with most efforts focusing on the text modality. ChatHaruhi [23] introduced a retrieval-augmented framework for character-consistent dialogue, while RoleLLM [36], Character-LLM [32], and RoleFact [31] proposed benchmarks and profile-grounded generation methods to improve persona simulation and mitigate hallucinations. Beyond text, MMRole [11] extended role-playing to the visual domain with a multimodal training and evaluation framework, and OmniCharacter [46] demonstrated the value of speech for real-time, voice-aware character interaction.
To assess role-playing capabilities, benchmarks such as RoleEval [33], RoleInteract [6], SocialBench [7], and CharacterEval [34] evaluated behavioral consistency and persona fidelity, while INCHARACTER [39] and ECHO [24] introduced human-aligned, psychometric evaluation paradigms. These advances underscore the importance of role-playing as a core dimension of LLM alignment and interactivity. However, existing efforts primarily target text-based or visual modalities, leaving a significant gap in the systematic evaluation of speech-based role-playing, where vocal expression is crucial for realism, emotional depth, and user engagement.
Interactive Speech Agents.
In recent years, speech-interactive LLMs have emerged as a central topic in human-computer interaction, with mainstream approaches typically falling into two paradigms: cascaded and end-to-end. Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS). This modularity has led to their wide adoption in systems such as ChatGPT and AudioGPT [21]. Modern TTS models like F5TTS [10], FireRedTTS [19], E2-TTS [16], Spark-TTS [38], CosyVoice2 [14], MaskGCT [40], and XTTS [4] have significantly improved output quality, expressiveness, and controllability. However, cascaded pipelines often suffer from error propagation and limited cross-modal context integration.
End-to-end systems aim to unify speech perception and generation within a single model. Early efforts like SpeechGPT [45] incorporated discrete speech units into LLM vocabularies, while later models such as AudioPaLM [29] utilized vocoder-based representations for continuous speech modeling. Recent advances, including Moshi [13], LLaMA-Omni [17], and Qwen2.5-Omni [41], have pushed the boundaries further with dual-modality training, streaming support, and low-latency interactive response, enabling more natural, expressive, and efficient spoken interactions.
3 SpeechRole-Data
We introduce SpeechRole-Data, a large-scale dataset for speech-based role-play, featuring 98 roles and over 112k dialogues with reference audio and rich character profiles. To ensure personality and voice consistency, we follow a structured pipeline spanning role selection, dialogue generation, and audio processing. Details are in the appendix.
3.1 Text Data Construction
Role Selection.
Role Metadata Extraction.
For each character, we extract structured metadata to build a comprehensive profile, including (1) Role Profile (temperament, preferences), (2) Background (social identity, relationships), and (3) Character Lines (dialogues, monologues). This metadata forms the basis for generating personality-consistent dialogues.
Dialogue Generation.
Using the extracted metadata as a guide, we employ gpt-4.1-2025-04-14 [25] to generate conversations. For each of the 98 roles, we produce approximately 800 single-turn and 800 multi-turn dialogues, creating a substantial initial pool of text data.
Dialogue Filtering and Deduplication.
The generated data undergoes a multi-stage filtering process. First, we ensure contextual relevance by computing semantic similarity between dialogues and their source metadata, using all_MiniLM_L6_v2 [28] for English and text2vec-bge-large-chinese [42] for Chinese, discarding samples with scores below 0.4 for English and 0.45 for Chinese. To eliminate redundancy, we remove near-duplicate dialogues using similarity thresholds of 0.9 for English and 0.85 for Chinese.
| Splits | Characters | Samples (single-turn/multi-turn) | Speech Hours (user/character) |
| Train | 78 | 89,461 (43,082/46,379) | 875.25 (204.45/670.80) |
| Dev | 20 | 21,993 (10,627/11,366) | 203.60 (49.84/153.76) |
| Test | 98 | 392 (196/196) | 4.20 (0.94/3.26) |
3.2 Role Voice Collection
Voice Collection and Preprocessing.
To acquire authentic reference voices, we collect audio from sources like the game Genshin Impact (with permission) and other audiovisual works. All audio is anonymized and fragmented into brief utterances to comply with fair use. Raw streams are extracted using ffmpeg [18] and converted to mono 16 kHz WAV format.
Voice Cleaning and Annotation.
We use the open-source Emilia framework [20] to process raw audio, including source separation, speaker diarization, and voice activity detection. This yields clean 3–10 second single-speaker clips, transcribed by whisper-medium-v3 [26] and scored with DNSMOS P.835 OVRL [27].
Speaker Identification.
The diarization process outputs speaker indices without explicit identity information. To attribute the audio to the correct characters, we use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [12]) to analyze transcribed text. The models’ integrated results are then manually verified.
Reference Audio Ranking and Selection.
To select the most representative audio sample for each role, we rank all of their voice clips based on vocal similarity. We use the CAM++ model [35] to extract a speaker embedding for each clip. The clip with the highest average cosine similarity to all other clips from the same character is selected as the final reference audio, as it best captures the character’s quintessential vocal profile.
3.3 Dataset Statistics
Dataset Overview.
SpeechRole-Data comprises 98 distinct roles and 112k dialogue samples, including both single-turn and multi-turn conversations (see Table 1). Each role is accompanied by a detailed role profile and a reference audio clip that demonstrates its unique speaking style, such as timbre, rhythm, and prosody. Each dialogue sample contains both text and speech, along with contextual background information relevant to the role, thereby supporting realistic, evidence-based role-playing. During the evaluation phase, the test set consists of two single-turn and two multi-turn dialogues sampled for each role. The remaining data are divided into a training set (78 roles) and a validation set (20 roles). Detailed examples are provided in Figure 2.
Speech Duration.
We analyze the speech duration of users and roles. As shown in Figure 3, the duration of role utterances is generally longer, typically ranging from 10 to 20 seconds. This is mainly due to the narrative or expressive nature of their content. In contrast, user utterances are concentrated within 3 to 6 seconds, reflecting their concise and goal-oriented characteristics.
4 SpeechRole-Eval
To systematically assess SRPAs, we introduce SpeechRole-Eval, a comprehensive benchmark that evaluates model performance across three key dimensions: fundamental interaction ability, speech expressiveness, and role-playing fidelity. In contrast to prior work that primarily focuses on text, SpeechRole-Eval places special emphasis on the speech modality as a critical component of evaluation. The scoring process is fully automated, powered by a strong multimodal model, and has been validated against human judgments to ensure reliability. This section describes the task setup, evaluation metrics, and the overall evaluation pipeline.
4.1 Task Types
To assess the cognitive, affective, and social competencies of SRPAs, user prompts are categorized into three major task categories based on communicative intent and interactional focus: (1) Internal Reasoning, which concerns the character’s rich subjective world, including personal perspectives, emotional responses, and underlying motivations; (2) Experiential Narration, which emphasizes the character’s ability to vividly recall past events, describe concrete actions, and provide relevant background or historical information; and (3) Social Communication, which highlights interpersonal understanding achieved through nuanced evaluative judgments, the provision of advice, and relational reasoning in context. These categories reflect the core dimensions of role-playing dialogue and enable a comprehensive, structured assessment of role-playing abilities (see Figure 2).
4.2 Evaluation Metrics
In contrast to traditional text-based RPAs, SRPAs must not only accurately emulate specific characters but also consistently deliver coherent and expressive spoken responses fully aligned with both the dialogue context and the character’s unique persona. To comprehensively evaluate SRPAs, we propose a robust three-dimensional evaluation framework that assesses fundamental interaction ability, speech quality and expressiveness, and role-playing fidelity.
Fundamental interaction abilities comprehensively evaluate the agent’s capacity to maintain coherent, natural, and instruction-aligned interactions fully consistent with the character, which are assessed by the following key metrics:
Instruction Adherence (IA): Do the spoken responses strictly follow the task instruction, remaining fully in character without any out-of-role explanations or assistant-like meta-comments?
Speech Fluency (SF): Are the responses delivered fluently, with smooth articulation, appropriate pacing, and minimal disfluencies such as stuttering or unnatural pauses?
Conversational Coherence (CC): Do the responses maintain logical consistency within the dialogue, aligning with previous content without contradictions or abrupt topic shifts?
Speech quality and expressiveness measure the naturalness and prosodic expressiveness of the generated speech, which are assessed by the following metrics:
Speech Naturalness (SN): Do the responses sound natural, human-like, and free from noticeable artifacts or robotic effects typically associated with synthetic speech?
Prosodic Consistency (PC): Does the prosody, including pitch, stress, and intonation, align with the character’s intended speaking style and remain consistent across the discourse?
Emotion Appropriateness (EA): Are emotional cues in the speech (e.g., anger, joy, sadness) well-aligned with the dialogue context and the character’s emotional state?
Role-playing fidelity assesses the extent to which the agent embodies the assigned character’s persona, which is assessed by the following metrics:
Personality Consistency (PeC): Do the responses consistently reflect the character’s personality traits, such as optimism, sarcasm, or authority?
Knowledge Consistency (KC): Are the responses grounded in the character’s established background, knowledge, and relationships, without fabricating out-of-character facts?
4.3 Evaluation Procedure
LLM-Based Scoring Procedure.
There is a lack of widely recognized objective metrics for the evaluation of role speech output, and direct scoring by LLMs often exhibits instability and inconsistency [37]. Inspired by Dai et al. [11], two strategies are adopted to enhance the reliability of LLM-based scoring.
First, the model is required to generate a brief rationale for its score prior to providing the original rating. This process encourages more deliberate and interpretable evaluations by the model. Second, rather than evaluating individual speech samples in isolation, a paired comparison setting is designed. For each evaluation instance, the model receives a test speech sample and a high-quality reference answer simultaneously, and assigns scores to both.
The final evaluation score is calculated as the ratio of the test speech score to the reference score, which standardizes the results across different prompts and reduces variance caused by prompt discrepancies. This two-step, reference-based evaluation framework is designed to improve scoring stability and promote better alignment with human judgments. We use gemini-2.5-pro-preview-06-05 [3] to assign scores to each response according to the predetermined evaluation protocol, with scores ranging from 1 to 10.
Human Correlation Verification.
To verify the reliability of our automatic scoring method, we randomly select 20 dialogue samples and, for each sample, randomly choose two model-generated responses. Four independent human annotators conduct pairwise comparisons of these responses based on multiple evaluation metrics, determining which response performs better for each metric. The final human score for each metric is calculated as the average of all annotators’ judgments. The results show a strong correlation between human evaluation and automatic scoring, confirming that our scoring method aligns closely with human judgments and provides a reliable foundation for large-scale automated evaluation. The human judgment setting follows the evaluation protocol introduced by Dai et al. [11], which uses pairwise preference comparisons to ensure stable and interpretable annotations. Details are in the appendix.
| Models | Evaluation Metrics | Overall | |||||||
| IA | SF | CC | SN | PC | EA | PeC | KC | ||
| English Evaluation Results | |||||||||
| Alibaba Cloud API | 0.9012 | 0.6678 | 1.0320 | 0.6973 | 0.6651 | 0.8087 | 0.8594 | 0.9171 | 0.8186 |
| XTTS | 0.9617 | 0.5493 | 0.9872 | 0.6578 | 0.7568 | 1.0133 | 1.0357 | 0.9531 | 0.8644 |
| CosyVoice2 | 1.0102 | 0.7869 | 1.0346 | 0.8682 | 0.9857 | 1.0781 | 1.0476 | 0.9744 | 0.9732 |
| E2 TTS | 1.0531 | 0.7938 | 1.1621 | 0.8859 | 1.0718 | 1.1806 | 1.1760 | 1.0104 | 1.0417 |
| F5-TTS | 1.1245 | 0.8057 | 1.1132 | 0.8717 | 0.9773 | 1.2686 | 1.0952 | 1.0614 | 1.0397 |
| FireRedTTS | 0.9587 | 0.7266 | 1.0297 | 0.8135 | 0.8760 | 1.0557 | 1.0415 | 0.9899 | 0.9365 |
| Spark-TTS | 1.0384 | 0.7583 | 1.0641 | 0.8282 | 1.0554 | 1.1427 | 1.0741 | 1.0144 | 0.9970 |
| GPT-SoVITS | 0.8566 | 0.4783 | 0.9293 | 0.5576 | 0.5606 | 0.8253 | 0.8899 | 0.8770 | 0.7468 |
| MaskGCT | 0.9277 | 0.7635 | 0.9902 | 0.7589 | 0.9449 | 1.0382 | 0.9701 | 1.0098 | 0.9254 |
[…]
| Chinese Evaluation Results | |||||||||
| Alibaba Cloud API | 0.8988 | 0.8473 | 0.8942 | 0.9275 | 1.0223 | 1.0493 | 0.9124 | 0.8887 | 0.9301 |
| XTTS | 0.6563 | 0.4951 | 0.7149 | 0.5348 | 0.4681 | 0.4945 | 0.6191 | 0.6595 | 0.5803 |
| CosyVoice2 | 0.7600 | 0.7739 | 0.7190 | 0.7943 | 0.7388 | 0.8843 | 0.6587 | 0.6597 | 0.7486 |
| E2 TTS | 0.6866 | 0.6513 | 0.7171 | 0.7082 | 0.6776 | 0.7027 | 0.6787 | 0.7161 | 0.6923 |
| F5-TTS | 0.7285 | 0.6901 | 0.7658 | 0.8052 | 0.8392 | 0.9291 | 0.6969 | 0.6890 | 0.7680 |
| FireRedTTS | 0.7536 | 0.6932 | 0.6968 | 0.7221 | 0.8371 | 0.8235 | 0.7093 | 0.6516 | 0.7359 |
| Spark-TTS | 0.7591 | 0.7179 | 0.7869 | 0.7887 | 0.8930 | 0.7902 | 0.7268 | 0.6890 | 0.7690 |
| GPT-SoVITS | 0.6915 | 0.6345 | 0.7269 | 0.6119 | 0.6583 | 0.6515 | 0.6296 | 0.6277 | 0.6540 |
| MaskGCT | 0.7391 | 0.5913 | 0.6558 | 0.7349 | 0.8011 | 0.8109 | 0.7722 | 0.6433 | 0.7186 |
[…]
5 Experiments
This section details our experimental setup and main results. We evaluate SRPAs under two paradigms: cascaded systems, which separate role-playing into ASR, LLM-based reasoning, and TTS stages, and end-to-end models that generate speech directly from prompts. All models are tested on the SpeechRole-Eval benchmark (Section 4) covering eight core metrics, with speaker similarity assessed for voice-controllable models. We also include supplementary experiments, such as fine-tuning a SRPA on our dataset to assess its speech role-playing performance.
5.1 SRPA Architectures
The primary objective of SRPAs is to generate coherent, spoken responses that are consistent with the target character’s identity upon receiving a user’s spoken query. To explore this, we examine two representative system architectures: the cascaded and end-to-end approaches.
Cascaded SRPAs decompose the overall task into three modular components. First, an ASR model transcribes the user’s spoken input into text. Next, a LLM generates a contextually appropriate and character-specific text-based response that aligns with the specified character’s identity, using both the transcription and system prompts defining the role. Finally, a TTS module converts the generated text into speech that matches the target character’s voice. In our implementation, both proprietary and open-source cascaded systems are explored. We utilize a proprietary cascaded system based on the Alibaba Cloud API [1], integrating Paraformer Realtime ASR v2, Qwen-Plus-Character, and CosyVoice-v2. For the open-source side, we employ whisper-large-v3-turbo [26] for high-quality ASR, LLaMA3.1-8B [15] for flexible and powerful text generation, and multiple state-of-the-art zero-shot timbre cloning TTS models for natural and expressive speech synthesis, including XTTS [4], CosyVoice2 [14], E2 TTS [16], F5-TTS [10], FireRedTTS [19], GPT-SoVITS [30], Spark-TTS [38], and MaskGCT [40].
5.2 Evaluation Results
Main Evaluation Results.
We report the main evaluation results of all SRPAs on the SpeechRole-Eval benchmark. This benchmark covers eight core metrics: Instruction Adherence (IA), Speech Fluency (SF), Conversational Coherence (CC), Speech Naturalness (SN), Prosody Consistency (PC), Emotion Appropriateness (EA), Personality Consistency (PeC), and Knowledge Consistency (KC), with the final score computed as the average across these dimensions.
As shown in Table 2, the cascaded agents generally outperform end-to-end systems across most evaluation dimensions. Notably, models such as E2 TTS, F5-TTS, and Spark-TTS achieve relatively high overall scores in both content relevance and expressiveness, benefiting greatly from their modular design and optimization. While end-to-end models, particularly GPT-4o Audio, demonstrate competitive performance in expressiveness and coherence, they still exhibit noticeable shortcomings in style consistency and character setting. The performance of Qwen2.5-Omni is significantly lower across all metrics, indicating that end-to-end modeling remains challenging in the absence of sufficient task-specific supervision.
Language-Level Observations.
We observe significant differences between the English and Chinese subsets. Overall, the model performs better on English, particularly with respect to metrics such as content coherence and knowledge consistency. This disparity likely reflects the model’s language imbalance in pre-training and TTS capabilities. For example, XTTS performs relatively well in English but shows a marked decline in Chinese, especially in the expression of emotion and character personality. In contrast, Alibaba Cloud API and CosyVoice2 demonstrate more stable performance across both languages. These results indicate that current SRPAs exhibit uneven performance across different languages, with significantly better results in English than in Chinese. Therefore, further efforts are necessary to develop high-quality multilingual SRPAs.
Speaker Similarity Comparison.
We further analyze the performance of SRPAs with voice cloning capability in terms of speaker similarity. Since end-to-end models do not support speaker cloning, we evaluate speaker similarity only for cascade systems. Following the same evaluation protocol as Seed-TTS [2], we use the WavLM-large [9] model to extract speaker embeddings and compute cosine similarity scores.
As shown in Figure 4, in the English setting, E2 TTS and F5-TTS achieve the highest speaker similarity, while GPT-SoVITS performs the worst. The overall ranking of models remains largely consistent across the two languages, indicating stable voice cloning performance. Notably, MaskGCT slightly outperforms others in Chinese, reaching the highest SIM score.
| Models | Evaluation Metrics | Overall | |||||||
| IA | SF | CC | SN | PC | EA | PeC | KC | ||
| English Evaluation Results | |||||||||
| Qwen2.5-Omni | 0.3977 | 0.4968 | 0.5491 | 0.4505 | 0.2958 | 0.3814 | 0.4089 | 0.4917 | 0.4340 |
| SpeechRole-Agent | 0.8072 | 0.6790 | 0.9480 | 0.6123 | 0.4202 | 0.5383 | 0.7568 | 0.9169 | 0.7098 |
[…]
| Chinese Evaluation Results | |||||||||
| Qwen2.5-Omni | 0.4924 | 0.6886 | 0.5832 | 0.6204 | 0.4693 | 0.5125 | 0.3932 | 0.4772 | 0.5296 |
| SpeechRole-Agent | 0.9188 | 0.8233 | 0.9037 | 0.7749 | 0.7277 | 0.8607 | 0.8402 | 0.9253 | 0.8468 |
[…]
5.3 Fine-Tuning SRPA
Training Setup.
We fine-tune Qwen2.5-Omni-7B on the SpeechRole training set, containing 78 roles with speech-based dialogues. The model trains using speech inputs paired with role-consistent responses, enabling it to better understand and generate role-based spoken interactions. During evaluation, we use SpeechRole-Eval, dividing it into two subsets: in-domain, which includes roles seen during training, and out-of-domain, consisting of unseen roles. Results are reported separately for English and Chinese.
Performance Analysis.
After fine-tuning on the complete SpeechRole-Data training set, SpeechRole-Agent achieves notable gains over the original Qwen2.5-Omni. In the in-domain test set, the overall score in English increases from 0.4340 to 0.7098, while in Chinese it rises from 0.5296 to 0.8468. All evaluation metrics improve, with particularly notable advancements in instruction adherence, emotional appropriateness, and personality consistency. In the out-of-domain test set, the model also demonstrates strong generalization ability, with the English score increasing from 0.4786 to 0.6795 and the Chinese score from 0.5117 to 0.8028. These results fully demonstrate the model’s robust role-playing capabilities beyond the scope of the training set. Detailed results are presented in Table 3.
6 Conclusion
We propose SpeechRole, a unified framework for developing and evaluating SRPAs. To support this objective, we construct SpeechRole-Data, a large-scale, high-quality dataset comprising 112k dialogues from 98 characters, each with rich role settings and diverse vocal characteristics. Our benchmark establishes rigorous evaluation protocols for role-consistent speech generation and introduces SpeechRole-Eval, an automated multimodal evaluation suite that assesses various aspects including interaction competence, speech expressiveness, and role consistency.
Experimental results indicate current SRPAs still face challenges in achieving fully coherent and expressive role simulation, underscoring the need for more advanced character-based speech behavior modeling. We expect this work will provide a solid foundation for future research at the intersection of speech generation, role-playing, and embodied agents. All data, evaluation protocols, and baseline methods will be made publicly available to facilitate further studies on speech-interactive agents.
Appendix A SpeechRole-Data Construction
A.1 Text Data Construction
Role Selection.
A total of 98 representative roles are meticulously curated for SpeechRole-Data, selected from two high-quality datasets: ChatHaruhi [23] and RoleLLM [36]. Specifically, 18 roles from games, movies, and television series are chosen from ChatHaruhi, while 80 roles from animation, movies, and television series are drawn from RoleLLM. The selected roles are balanced in terms of language, gender, and personality traits, thereby ensuring diversity in both vocal expression and character identity (see Figure 5). To facilitate rich and character-consistent speech modeling, particular emphasis is placed on roles with distinctive vocal characteristics. Table 5 and Table 6 summarizes the attributes of all roles, including their categories, languages, genders, source materials, and dataset splits.
Role Metadata Extraction.
For each character, we extract structured metadata, including: (1) Role Profile, such as temperament, behavioral tendencies, values, and personal preferences; (2) Background, including character settings, social identity, relationships, and living environment; and (3) Character Lines, including narration, monologues, and dialogue excerpts involving the target character in games or audiovisual works. These elements collectively define distinctive character profiles and support vivid, context-rich speech role-playing. Additionally, the dialogue excerpts serve as the basis for guiding dialogue generation and evaluating character consistency.
Dialogue Generation.
We use gpt-4.1-2025-04-14 [25] to generate dialogues between users and roles based on role metadata. For some roles with lengthy scripts, the content is divided into segments and processed sequentially. GPT-4.1 is required to provide relevant script segments while generating dialogues to serve as references for SRPA in answering questions. For each role, 800 single-turn dialogues and 800 multi-turn dialogue data are generated.
Dialogue Filtering.
The data are filtered through a multi-stage process to ensure structural consistency, contextual relevance, and semantic diversity. First, only dialogues that conform to a clear turn-taking pattern are retained. In these dialogues, the user initiates with a question, and the role responds, with each party speaking in alternating turns. To ensure alignment between the dialogue and its corresponding context, semantic embeddings are computed for both texts. For English data, the all_MiniLM_L6_v2 [28] model is used, and samples with a similarity score below 0.4 are excluded. The same procedure is applied to Chinese data using the text2vec-bge-large-chinese [42] model, with a similarity score threshold of 0.45.
To further eliminate semantic redundancy, which is commonly observed in GPT-4.1-generated dialogues, near-duplicate samples are removed based on the similarity of their dialogue embeddings. Thresholds are set to 0.85 for Chinese and 0.9 for English. This de-duplication step is conducted separately for single-turn and multi-turn dialogues for each role. After filtering, the dataset contains 53,902 single-turn dialogues and 57,944 multi-turn dialogues.
| Metrics | IA | SF | CC | SN | PC | EA | PeC | KC | Overall |
| MSE | 0.1020 | 0.0413 | 0.1426 | 0.0581 | 0.2539 | 0.1262 | 0.1690 | 0.1147 | 0.1260 |
| Pearson | 0.7304 | 0.6366 | 0.6076 | 0.8093 | 0.6316 | 0.8881 | 0.6677 | 0.7581 | 0.7162 |
A.2 Role Voice Collection
Role Voice Extraction.
To achieve authentic voice role-playing, real-life performance audio for each character is collected as reference material. For characters from Genshin Impact, in-game character audio is collected with permission from miHoYo. For other characters, audio is obtained from actual film and television works. Films and television series featuring these characters are gathered, and a total of 105 video files with an approximate duration of 181 hours are downloaded from the internet, covering all target characters.
During the preprocessing stage, all extracted speech segments undergo anonymization and fragmentation to ensure that they contain only brief, decontextualized utterances, with no retention of the original video or narrative context. As a result, no identifiable or copyrighted content remains, thus mitigating potential intellectual property risks and ensuring adherence to the principles of fair use for research purposes.
We first employ a batch processing pipeline based on ffmpeg [18] to extract the raw audio streams from the collected video files. To ensure consistency, all audio is converted to the standard mono 16 kHz WAV format. However, since these recordings often contain background music, noise, and overlapping speech, additional processing is required.
Role Voice Processing.
To obtain clean audio segments from specific speakers, we employ a structured pipeline based on the open-source Emilia [20] framework. This process includes source separation to remove background noise, speaker separation to distinguish individual speakers, and voice activity detection for fine-grained segmentation. The resulting audio segments, each lasting between 3 and 10 seconds and containing only a single speaker, are well-suited for subsequent modeling.
Finally, we use an automatic speech recognition model to transcribe each segmented audio fragment and evaluate speech quality using the DNSMOS P.835 OVRL [27] metric. In order to ensure the quality of speech data, we filter out speech with DNMOS less than 3. The resulting structured dataset provides, for each audio entry, the following information: start and end times, transcription, speaker index, language, and DNSMOS-based audio quality score.
Speaker Identification.
The extracted audio data contains only the speaker’s index and does not include identity information. To identify the speaker, we use both gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [12] models to determine the target character’s speaker number based on all transcribed texts. We integrate the results from both models and verify them manually to obtain the speech corresponding to the target character.
Role Voice Ranking.
To select the most representative audio sample as a reference for subsequent role-playing, we rank all speech samples of each speaker based on similarity. Specifically, we extract the speaker feature vector of each audio sample using the CAM++ [35] model and calculate the cosine similarity between each sample and all other samples. A higher average similarity score indicates that the audio sample better represents the overall speaker characteristics of the role. Finally, we sort all audio samples for each role in descending order according to their similarity scores and select the top-ranked sample as the reference audio.
Appendix B Human Correlation Verification Details
To evaluate the consistency between automatic scoring and human scoring, we randomly selected 20 samples from the dataset, including 5 single-turn dialogue samples and 15 multi-turn dialogue samples. For each sample, we randomly selected two model-generated speech responses for evaluation.
Gemini scored each model response and its corresponding reference response on a scale of 1 to 10 based on 8 predefined scoring metrics. The final score for a model was calculated as the ratio of the model response score to the reference response score.
For human evaluation, four trained evaluators were employed to directly compare the responses of the two models based on the same 8 scoring metrics. The evaluation criteria were as follows: the superior model received 1 point, the inferior model 0 points, and if the performance was deemed equivalent, each received 0.5 points. The final human score was determined by averaging the scores from all evaluators.
To quantify the agreement between automated and human evaluation methods, we computed two key metrics: mean squared error (MSE) and Pearson correlation coefficient. The MSE was calculated by first fitting a linear regression between the score differences of paired model responses from human evaluators and gemini. The predicted gemini score differences were then derived from this regression model, and the MSE was computed between these predicted values and the actual gemini score differences. For the Pearson correlation analysis, we computed the linear correlation coefficient between the human evaluators’ score differences and gemini’s score differences across all paired comparisons.
Table 4 reports the detailed results, the overall MSE of 0.1260 indicates relatively minor deviations between the two scoring methods. The high Pearson correlation coefficient further confirms the reliability of the automated scoring system. The evaluation results collectively demonstrate strong agreement between automatic and human scoring across all metrics, indicating that the automated scoring system can effectively approximate human judgment.
Appendix C Additional data statistics
Role Type Distribution
Figure 5 presents the distribution of characters by language (Chinese and English), source type (game, film, television series, and animation), and gender. English characters, particularly those from films, constitute the majority in the dataset. In contrast, the Chinese subset exhibits a more balanced distribution across various source types. This diversity provides extensive coverage of voices and contexts for role-playing.
Task Type Distribution
User prompts can generally be categorized into three major task categories, which can be further divided into eight subtask types, as illustrated in the figure 5. The description and proportion of each subtask type are as follows:
1.
Internal Reasoning
Opinion and Emotion Inquiry: Questions about opinions, emotions, inner feelings, or reactions. (14.73%)
Reasoning and Motivation: Questions about why someone did something, their motivations, or decision-making processes. (14.73%)
Reflection and Change: Questions about how practices/opinions have changed. (9.55%)
2.
Experiential Narration
Action Description and Retrospection: Questions about what happened, how something was done, or detailed retrospectives. (14.72%)
Skills and Abilities: Questions about how to accomplish something, difficulties with skills, or ability changes. (6.38%)
Historical and Background Description: Questions about history or background of places/organizations/people. (11.55%)
3.
Social Communication
Judgment of Others and Events: Questions about relationships, event evaluations, or interpersonal interactions. (14.73%)
Advice and Life Experience: Questions seeking advice, coping methods, or life experiences. (13.61%)
Evaluation Results Demonstration
Figure 6 visually presents the evaluation results of all SRPAs on the SpeechRole-Eval benchmark.
Appendix D Prompt Templates
Prompts for Batch Data Generation
Prompts for Automated Judgement
Figure 11 displays the judgment prompt designed for automatically assessing audio responses generated by the two models.
Appendix E SpeechRole-Agent training settings
Figure 12 presents the distributed training script for fine-tuning the Qwen2.5-Omni-7B model on 8×H20 GPUs. The configuration employs bfloat16 mixed-precision training with a learning rate of 1e-4 and gradient accumulation. The implementation includes periodic evaluation (every 500 steps) and checkpoint management.
| Role Name | Category | Language | Gender | Source | Split |
| hutao | Game | Chinese | Female | Genshin Impact | train/test |
| raidenShogun | Game | Chinese | Female | Genshin Impact | train/test |
| wanderer | Game | Chinese | Male | Genshin Impact | train/test |
| ayaka | Game | Chinese | Female | Genshin Impact | dev/test |
| zhongli | Game | Chinese | Male | Genshin Impact | train/test |
| liyunlong | TV | Chinese | Male | Drawing Sword | dev/test |
| wangduoyu | Movie | Chinese | Male | Hello Mr. Billionaire | train/test |
| weixiaobao | TV | Chinese | Male | The Deer and the Cauldron | train/test |
| jiumozhi | TV | Chinese | Male | Demi-Gods and Semi-Devils | train/test |
| wangyuyan | TV | Chinese | Female | Demi-Gods and Semi-Devils | train/test |
| Luna | Movie | English | Female | Harry Potter | dev/test |
| Penny | TV | English | Female | The Big Bang Theory | dev/test |
| zhangwuji | TV | Chinese | Male | The Heaven Sword and Dragon Saber | train/test |
| zhaomin | TV | Chinese | Female | The Heaven Sword and Dragon Saber | train/test |
| huangrong | TV | Chinese | Female | The Legend of the Condor Heroes | train/test |
| guojing | TV | Chinese | Male | The Legend of the Condor Heroes | dev/test |
| wukong | TV | Chinese | Male | Journey to the West | train/test |
| HAL 9000 | Movie | English | Male | 2001: A Space Odyssey | train/test |
| Colonel Nathan R. Jessep | Movie | English | Male | A Few Good Men | train/test |
| Antonio Salieri | Movie | English | Male | Amadeus | train/test |
| Stifler | Movie | English | Male | American Pie | train/test |
| Paul Vitti | Movie | English | Male | Analyze That | train/test |
| Alvy Singer | Movie | English | Male | Annie Hall | train/test |
| Violet Weston | Movie | English | Female | August: Osage County | train/test |
| Willie Soke | Movie | English | Male | Bad Santa | train/test |
| Gaston | Animation | English | Male | Beauty and the Beast | train/test |
| The Dude | Movie | English | Male | The Big Lebowski | train/test |
| Paul Conroy | Movie | English | Male | Buried | train/test |
| Truman Capote | Movie | English | Male | Capote | train/test |
| Mater | Animation | English | Male | Cars 2 | train/test |
| Andrew Detmer | Movie | English | Male | Chronicle | train/test |
| Coriolanus | Movie | English | Male | Coriolanus | train/test |
| John Keating | Movie | English | Male | Dead Poets Society | dev/test |
| Wade Wilson | Movie | English | Male | Deadpool | dev/test |
| Jim Morrison | Movie | English | Male | The Doors | train/test |
| Queen Elizabeth I | Movie | English | Female | Elizabeth: The Golden Age | dev/test |
| Jeff Spicoli | Movie | English | Male | Fast Times at Ridgemont High | train/test |
| Fred Flintstone | Animation | English | Male | The Flintstones | train/test |
| Freddy Krueger | Movie | English | Male | Freddy Vs.Jason | train/test |
| Tyrion Lannister | TV | English | Male | Game of Thrones | train/test |
| James Brown | Movie | English | Male | Get on Up | train/test |
| Walt Kowalski | Movie | English | Male | Gran Torino | train/test |
| John Coffey | Movie | English | Male | The Green Mile | train/test |
| Theodore Twombly | Movie | English | Male | Her | dev/test |
| Gregory House | TV | English | Male | House M.D. | dev/test |
| Sonny | Movie | English | Male | I, Robot | train/test |
| Colonel Hans Landa | Movie | English | Male | Inglourious Basterds | train/test |
| Judge Dredd | Movie | English | Male | Judge Dredd | dev/test |
| Juno MacGuff | Movie | English | Female | Juno | train/test |
| Professor G.H. Dorr | Movie | English | Male | The Ladykillers | train/test |
| Role Name | Category | Language | Gender | Source | Split |
| Fletcher Reede | Movie | English | Male | Liar Liar | train/test |
| Abraham Lincoln | Movie | English | Male | Lincoln | train/test |
| Frank T.J. Mackey | Movie | English | Male | Magnolia | train/test |
| Leonard Shelby | Movie | English | Male | Memento | train/test |
| Harvey Milk | Movie | English | Male | Milk | train/test |
| Randle McMurphy | Movie | English | Male | One Flew Over the Cuckoo’s Nest | train/test |
| Jack Sparrow | Movie | English | Male | Pirates of the Caribbean: Dead Man’s Chest | dev/test |
| John Dillinger | Movie | English | Male | Public Enemies | train/test |
| Lestat de Lioncourt | Movie | English | Male | The Queen of Damned | train/test |
| Tyler Hawkins | Movie | English | Male | Remember Me | dev/test |
| James Carter | Movie | English | Male | Rush Hour 2 | train/test |
| Jigsaw | Movie | English | Male | Saw | train/test |
| John Doe | Movie | English | Male | Se7en | train/test |
| Sherlock Holmes | Movie | English | Male | Sherlock Holmes | dev/test |
| Shrek | Animation | English | Male | Shrek | train/test |
| Pat Solitano | Movie | English | Male | Silver Linings Playbook | train/test |
| Karl Childers | Movie | English | Male | Sling Blade | train/test |
| Bruno Antony | Movie | English | Male | Strangers on a Train | train/test |
| Seth | Movie | English | Male | Superbad | train/test |
| Caden Cotard | Movie | English | Male | Synecdoche, New York | train/test |
| Travis Bickle | Movie | English | Male | Taxi Driver | train/test |
| Stanley Ipkiss | Movie | English | Male | The Mask | dev/test |
| Lyn Cassady | Movie | English | Male | The Men Who Stare at Goats | train/test |
| Michael Scott | TV | English | Male | The Office | dev/test |
| Robert Angier | Movie | English | Male | The Prestige | dev/test |
| Dr. Frank-N-Furter | Movie | English | Male | The Rocky Horror Picture Show | train/test |
| Jack Torrance | Movie | English | Male | The Shining | train/test |
| Tom Ripley | Movie | English | Male | The Talented Mr. Ripley | train/test |
| D_Artagnan | Movie | English | Male | The Three Musketeers | train/test |
| Thor | Movie | English | Male | Thor: Ragnarok | train/test |
| James Bond | Movie | English | Male | Tomorrow Never Dies | dev/test |
| Mark Renton | Movie | English | Male | Trainspotting | train/test |
| David Aames | Movie | English | Male | Vanilla Sky | train/test |
| Rorschach | Movie | English | Male | Watchmen | train/test |
| Jordan Belfort | Movie | English | Male | The Wolf of Wall Street | train/test |
| Logan | Movie | English | Male | X-Men Origins: Wolverine | dev/test |
| Judy Hoops | Animation | English | Female | Zootopia | train/test |
| Doctor Who | TV | English | Male | Doctor Who | train/test |
| Raylan Givens | TV | English | Male | Justified | train/test |
| Mary Sibley | TV | English | Female | Salem | train/test |
| Lucifer Morningstar | TV | English | Male | Lucifer | train/test |
| Twilight Sparkle | Animation | English | Female | My Little Pony: Friendship is Magic | dev/test |
| Oliver Queen | TV | English | Male | Arrow | train/test |
| Klaus Mikaelson | TV | English | Male | The Originals | train/test |
| Queen Catherine | TV | English | Female | Reign | train/test |
| Dr. Hannibal Lecter | TV | English | Male | Hannibal | train/test |
| Coach Eric Taylor | Movie | English | Male | Friday Night Lights | train/test |
| yaemiko | Game | Chinese | Female | Genshin Impact | train/test |